# Send Message to the Future? Blockchain-based Time Machines for Decentralized Reveal of Locked Information

### v2

@@ IV-A Applying the Blockchain-based Architecture
Figure 3 shows a high-level overview of the system with an example. To use the system, clients submit encrypted messages to the smart contract, and secret holders store the messages in encrypted form. The holders then decrypt and publish the messages on the blockchain at the specified decryption time determined by the client. Clients are not responsible for decrypting secret transactions, allowing them to exit the process after submitting the encrypted message. For example, a client can send an encrypted message at 3 o’clock and ask it to be committed at 4 o’clock.
@@
IV-B Addressing Design Challenges
The proposed architecture employs smart contracts to serve as the communication medium, providing a reliable communication channel and an incentive mechanism to ensure secret holders remain active and honest. According to the identified design challenges in the previous section, reliable condition check and verifiable reveal are two important features left to provide.
Reliable condition check
@@ IV-B Addressing Design Challenges
In addition to the centralized global clock, the system uses the power of the decentralized blockchain clock as an auxiliary time source. While distributed systems such as blockchains do not achieve perfect clock synchronization, they offer a resilient and available time reference with bounded drift given the safety and liveness of the blockchain [44, 45]. This blockchain clock is grounded in the concept of block time, which is manifested as timestamps embedded within block headers.
Although the blockchain clock may exhibit a lower degree of precision compared to the centralized global clock, it plays a crucial role of fail-safe mechanism in our system to ensure that secrets are not revealed earlier or later in time. For instance, the Ethereum network mandates validators to synchronize their time with the network upon joining [46, 47]. In cases where a validator’s local time deviates significantly from the network time, it results in isolated blocks that are not accepted by the majority of the network. Implementation wise, it only accommodates a 15-second forward time difference between a validator’s local clock and the clock of the block proposer [48], which means a malicious block proposer can only advance the Ethereum network’s time by a bounded 15-second window. This provides additional resistance against early message decryption. In the extreme case where a majority of secret holders’ local clock failed and an attacker is chosen as the block proposer involving message decryption, Ethereum still serves as an effective tool in preventing the 15-second onward advanced disclosure of the message. By incorporating these two complementary time references, the proposed system strikes a balance between precision and robustness, ensuring the secure and timely release of confidential information.
Verifiability of Revealed Information
@@ IV-C Example Workflow
Upon reaching the client-specified decryption time, honest holders proceed to submit their shares to the smart contract, signifying the reveal of the secret. Subsequently, holders retrieve and locally verify the shares submitted by peer holders. If no disputes are raised, the smart contract allocates rewards to the first three holders who successfully submit their shares. Despite submitting a correct share, the fourth holder does not receive a reward, as its contribution is redundant for the secret reveal. This reward strategy incentivizes timely submission of secret shares.
@@
V Fully Verifiable Secret Sharing on Public Communication Channel
A vital requirement in designing the proposed system is to provide verifiability to the correctness of message decryption. This section presents a new cryptographic protocol to fulfill this requirement. The proposed protocol achieves linear communication cost when the number of participant increases, without the need for a peer-to-peer communication channel. The proposed protocol is not only applicable to the timed release cryptography system, but also to other secret sharing scenarios, where only a public communication channel is available. The relevant math symbols are listed in Table II.
@@ V-A Objectives and Problem Modelling
The objective of the cryptographic protocol is to hide, distribute, and later recover a piece of information $k$. When applied to the timed-release cryptography system, $k$ is a symmetric key to encrypt and decrypt a time-sensitive message $m$. Two roles are involved in this protocol, on the one hand, it is the client that holds $k$, on the other hand there are multiple secret holders that receives shares of $k$ for recovery in the future. Since the proposed architecture is blockchain-based, it is required that clients and secret holders only communicate via the blockchain, which is actually a broadcasting channel.
@@ V-B Design
With secret shares communicated through Diffie-hellman, bilinear pairing can be adopted to achieve verifiability of the secret shares, i.e. $s_{i}$ revealed by holder $i$ is correct. Note that Diffie-Hellman key exchange is secure against attackers that aim to find out the exchanged keys in any fields where discrete logarithm is hard to compute, therefore, it is applicable in BLS12-381, a pairing-friendly curve, on which discrete logarithm is hard. Therefore $s_{i}$ should be computed by $s_{i}=g_{1}^{r\cdot sk_{i}}$ and the client should also broadcast $g_{1}^{r}$. On BLS12-381, $s_{i}$ can be verified by comparing $e(pk_{i},g_{2}^{r})$ and $e(s_{i},g_{2})$. This is correct because $e(pk_{i},g_{2}^{r})=e(g^{sk_{i}},g_{2}^{r})=e(g^{r\cdot sk_{i}},g_{2})$. To perform this check, a client only needs to publish $g_{2}^{r}$, regardless of the number of holders.
The next step is to establish a connection between these secret shares and $k$, such that any $t$ out of $n$ secret values can recover $k$. Similar to the existing secret sharing schemes, we hide $k$ in the evaluation of a $t-1$ degree polynomial. This polynomial $P$ is constructed by interpolating the points $(1,s_{1}),...,(t-1,s_{t-1})$ along with $(0,k)$. The remaining shares, $s_{t},...,s_{n}$ form a relation with $k$ by transforming them into $P(t),...,P(n)$. The transformation is achieved with encryption and decryption. Regard $P(t),...,P(n)$ as messages to encrypt, $s_{t},...,s_{n}$ are the symmetric keys to encrypt these messages, the ciphertexts denoted as $\alpha_{t},...,\alpha_{n}$ can be obtained and published. This means the owner of $s_{i}$, holder $i$, can use $s_{i}$ to decrypt $\alpha_{i}$ to obtain the evaluation of $P$ at $i$. Since $s_{i}$ is a one-time key, the encryption scheme is essentially a one-time pad. As a concrete example, the exclusive OR operation can be used as the encryption and decryption method, i.e. $\alpha_{i}=P(i)\oplus s_{i}$. Therefore, computing and publishing $\alpha_{t},...,\alpha_{n}$ allows holder $t$ to holder $n$ to obtain $P(t),...,P(n)$. To conclude, the secret shares $s_{i}$ gives each secret holder an distinct evaluation on $P$. For holder $1$ to holder $t-1$, $s_{i}$ equals to $P(i)$; for holder $t$ to holder $n$, $s_{i}$ is the key to decrypt the ciphertext $\alpha_{i}$ to obtain the plaintext $P(i)$.
@@
The verifiability of the recovered secret is provided by the verfiability of each individual secret share from bilinear pairing evaluation. Given that the secret shares are verifiabily correct and the $\alpha$s are public, anyone can use the same inputs on Lagrange interpolation to reproduce $k$.
@@ V-C Applying to Timed Release Cryptography
Combining all components in the design, we go over the entire process of the proposed cryptographic protocol when it is applied to the time timed release cryptography system. Suppose the client has a message $m$ to encrypt for a period of time; through the published smart contract that stores information of the system, the client knows the public keys of the secret holders $pk_{1},...,pk_{n}$.
@@
    When received a request from a client, get the secret share $s_{i}=pk_{i}^{r}=g_{1}^{r\cdot sk_{i}}$.
@@ VI-A Attack Model
Three types of adversaries are taken into account in the attack model, external adversaries, malicious secret holders, and malicious clients. External adversaries are malicious parties that aims to obtain confidential information using the public knowledge, specifically, the encrypted messages before decryption time.
@@
Malicious clients also have potential to disrupt the system if they can compromise the verifiability of decrypted messages. The system fails if clients can falsely accuse honest secret holders for submitting incorrect shares. In this case, honest secret holders are punished and disqualified even if they follow the protocol. Therefore, the proposed protocol must ensure verifiability of secret shares in the existence of malicious clients.
@@ VI-B Message Secrecy Before Decryption
For all generated and secure parameters of holders’ secret keys $sk$ and the clients’ choice of $r$, an external adversary can never use the public knowledge $(g_{1},g_{2},pk_{1},pk_{2},pk_{3},g_{1}^{r},g_{2}^{r},\alpha_{3})$ to derive a secret share before the corresponding holders submits it.
@@
With this base case of three secret holders. We extend this proof to show that message secrecy holds for any size of secret holders.
@@
We further provide the proof that the secrecy of the message key $k$ holds in the present of malicious secret holders below threshold.
@@
Suppose the collaborating malicious secret holders can obtain $k$ before honest holders submit shares, they must be able to gain one extra secret share belonging to the honest holders. That is, they are able to get the secret share using the public knowledge $g_{1},g_{2},g_{1}^{r},g_{2}^{r},pk_{1},...,pk_{n},\alpha_{t+1},...,\alpha_{n}$ plus their own secret keys. Since their own secret keys are randomly generated values, which are irrelevant to any other honest holders’ secret share, they must be able to obtain the secret share using the public knowledge. However, lemma VI.2 shows that it is not possible to obtain a secret share using the public knowledge, which contradicts the initial assumption and proves the lemma. ∎
@@ VII Performance Evaluation
The most fundamental requirement for a timed-release cryptography system is to decrypt messages at the time users specify. Therefore, decryption time deviation, computed as the difference between the expected decryption time and actual decryption time is measured in this section under various scenarios. Moreover, we provide an estimation of energy consumption of the proposed system, showing its advantage in energy efficiency.
@@ VII-A Decryption Time Stability
As indicated in Figure 7, the time lock puzzle approach exhibits unreliable decryption time. Its performance deteriorates with increasing encryption duration; for instance, a one-week encryption period results in an average decryption time deviation exceeding two hours. The result indicates two disadvantages of time lock puzzle systems. First, its the long error bars illustrate the unstable puzzle solving time. Second, the average lines at the early decryption area far from zero deviation show that the puzzle difficulty cannot be adjusted perfectly to achieve the desired average solving time. In the case of the Bitcoin blockchain, the hash rate of the network is in a strong increasing trend over time, while the increase of the puzzle difficulty does not strictly follow, resulting in an average block time shorter than expected, thus an accumulated deviation as the encryption duration increases.
@@ VII-B Scalability
In the case of the proposed system, the latency in the decryption process comprises two elements: the time needed for holders to submit secret shares to the blockchain and the time to verify a sufficient number secret shares. To understand the influence of scaling the number of secret holders on decryption latency, experiments of 3, 10, 20, 30, and 40 secret holders are conducted to process dummy time lock messages.
@@
In Figure 8, the publishing time of the secret shares shows the latency publishing a sufficient number of shares on the blockchain; the local verification time shows the latency of a secret holder getting enough secret shares verified. As holders process requests concurrently, increasing the number of holders does not extend the time required for the blockchain to accumulate a sufficient number of secret shares. However, as it requires more secret shares to validate a message when the number of secret holders scales, the time required for the secret holders to locally ensure the correct reconstruction of messages increases linearly.
@@ VIII Resilience to Malicious Voting
The sincere and malicious population of voters: In every simulation, a certain population of voters turn into malicious and change their votes based on the winners of the remaining population. So if $l$of $\mathcal{N}$ voters are sincere, $(1-l)$of $\mathcal{N}$ voters change their votes to jeopardize the decision outcomes of the $l$of the voters. We have a total of 100 simulations, where $l$ is varied from 1 to 100, with a increment of 1. In every simulation, we have 100 iterations, wherein the set of malicious voters ($(100-l)$%) changes due to random sampling.
@@
We consider real-world datasets for electing government leaders across different cities or countries (such as UK labour elections, elections in Dublin, Ireland, Oakland and Minneapolis elections in the United States [62, 63, 64, 65], where voters provide either complete or incomplete strict preferential order for the candidates. The election data are taken from the Preflib library [66], which hosts benchmarks for real-world voting and preference data.
We select the datasets to have a mixture of elections that have different numbers of candidates (varying from 5 to 9) or voters (varying from 266 to 36655). As demonstrated in Figure 9, all four elections suffer from a change of winner in the presence of malicious voters. For example, the UK Labour Leadership Election, a small-scale election with only 266 voters, shows a change of winner in the present of 40 strategic voters. While larger scale elections tolerate more malicious voters, change of winner is still possible with collaboration of potential malicious voters with similar preferences.
@@
In summary, a model is formulated to define malicious voting and apply it to real-world election datasets. This model is tested and validated on real-world datasets. We observe that if the purposed system is used, the probability of voters accessing or attacking the votes is negligible. Hence the proposed method is a secure solution to preserve fairness of real-world election outcomes.

### v3

@@ IV-A The Smart-Contract-based Architecture
Figure 3 shows a high-level overview of the system with an example. To use the system, clients submit encrypted messages to the smart contract, and secret holders store the messages in encrypted form. The holders then decrypt and publish the messages on the blockchain at the specified decryption time determined by the client. Clients are not responsible for decrypting secret messages, allowing them to exit the process after submitting the encrypted message. For example, a client can send an encrypted message at 3 o’clock and ask it to be committed at 4 o’clock.
@@
IV-B Improved Clock Reliability with On-Chain Timestamps
@@
In addition to the centralized global clock, the system uses the power of the decentralized blockchain clock as an auxiliary time source. While distributed systems such as blockchains do not achieve perfect clock synchronization, they offer a resilient and available time reference with bounded drift given the safety and liveness of the blockchain [50, 51, 52]. This blockchain clock is grounded in the concept of block time, which is manifested as timestamps embedded within block headers.
The blockchain clock plays a crucial role as a fail-safe mechanism in our system to ensure that secrets are not revealed earlier in time. Tzinas et al. [50] have shown that timestamps produced from blockchains have high accuracy. Take Ethereum as an example, the network mandates validators to synchronize their time with the network upon joining [53, 54]. In cases where a validator’s local time deviates significantly from the network time, it results in isolated blocks that are not accepted by the majority of the network. Implementation-wise, it only accommodates a 15-second forward time difference between a validator’s local clock and the clock of the block proposer [55], which means a malicious block proposer can only advance the Ethereum network’s time by a bounded 15-second window. This provides additional resistance against early message decryption. In the extreme case where a majority of secret holders’ local clock failed and an attacker is chosen as the block proposer involving message decryption, Ethereum still serves as an effective tool in preventing the 15-second onward advanced disclosure of the message. By incorporating these two complementary time references, the proposed system strikes a balance between precision and robustness, ensuring the secure and timely release of confidential information.
IV-C Verifiability of Revealed Information
@@ IV-D Example Workflow
Upon reaching the client-specified decryption time, honest holders proceed to submit their shares to the smart contract, signifying the reveal of the secret. Subsequently, holders retrieve and locally verify the shares submitted by peer holders. If no disputes are raised, the smart contract allocates rewards to the first three holders who successfully submit their shares. Despite submitting a correct share, the fourth holder does not receive a reward, as its contribution is redundant for the secret reveal. This reward strategy incentivizes timely submissions of secret shares.
@@
V Secret Sharing with Reveal-Verifiability on Public Communication Channel
The proposed protocol in this section is not only applicable to the timed-release cryptography system. This protocol itself makes an independent contribution to cryptographic secret sharing techniques by achieving reveal-verifiability and can be applied to other scenarios where verifiable message recovery is required.
A vital requirement in designing the proposed system is to provide verifiability to the correctness of message decryption. This section presents a new cryptographic protocol to fulfill this requirement. The proposed protocol achieves linear communication cost when the number of participants increases, without the need for a peer-to-peer communication channel. The relevant math symbols are listed in Table II.
@@ V-A Objectives and Problem Modelling
The objective of the cryptographic protocol is to hide, distribute, and later recover a piece of information $k$. When applied to the timed-release cryptography system, $k$ is a symmetric key to encrypt and decrypt a time-sensitive message $m$. Two roles are involved in this protocol, on the one hand, it is the client that holds $k$, on the other hand, multiple secret holders receive shares of $k$ for recovery in the future.
Unlike existing secret sharing protocols, since the proposed architecture is blockchain-based, it is required that clients and secret holders only communicate via sending permissionless blockchain transactions, i.e. in a publicly available broadcasting channel. Moreover, verifiability of revealed secret shares is an additional requirement compared to existing secret sharing methods.
We assume all honest parties are not compromised, i.e., secret holders keep their secret key safe, and clients keep their secret information $k$ safe. The broadcasting channel is assumed to be reliable such that all messages are eventually delivered to all parties. The order of the messages can be changed because each secret share is independent and can be revealed in arbitrary order.
@@ V-B Design
With secret shares communicated through Diffie-hellman, bilinear pairing can be adopted to achieve verifiability of the secret shares, i.e. $s_{i}$ revealed by holder $i$ is correct. Note that Diffie-Hellman key exchange is secure against attackers that aim to find out the exchanged keys in any fields where discrete logarithm is hard to compute, therefore, it is applicable in BLS12-381, a pairing-friendly curve, on which discrete logarithm is hard. Therefore $s_{i}$ should be computed by $s_{i}=g_{1}^{r\cdot sk_{i}}$ and the client should also broadcast $g_{1}^{r}$. On BLS12-381, $s_{i}$ can be verified by comparing $e(pk_{i},g_{2}^{r})$ and $e(s_{i},g_{2})$. This is correct because $e(pk_{i},g_{2}^{r})=e(g_{1}^{sk_{i}},g_{2}^{r})=e(g_{1}^{r\cdot sk_{i}},g_{2})$. To perform this check, a client only needs to publish $g_{2}^{r}$, regardless of the number of holders.
The next step is to establish a connection between these secret shares and $k$, such that any $t$ out of $n$ secret values can recover $k$. Similar to the existing secret sharing schemes, we hide $k$ in the evaluation of a $t-1$ degree polynomial. This polynomial $P$ is constructed by interpolating the points $(1,s_{1}),...,(t-1,s_{t-1})$ along with $(0,k)$. The remaining shares, $s_{t},...,s_{n}$ form a relation with $k$ by transforming them into $P(t),...,P(n)$. The transformation is achieved with encryption and decryption. Regard $P(t),...,P(n)$ as messages to encrypt, $s_{t},...,s_{n}$ are the symmetric keys to encrypt these messages, the ciphertexts denoted as $\alpha_{t},...,\alpha_{n}$ can be obtained and published. This means the owner of $s_{i}$, holder $i$, can use $s_{i}$ to decrypt $\alpha_{i}$ to obtain the evaluation of $P$ at $i$. Since $s_{i}$ is a one-time key, the encryption scheme is essentially a one-time pad. As a concrete example, the exclusive OR operation can be used as the encryption and decryption method, i.e. $\alpha_{i}=P(i)\oplus s_{i}$. Therefore, computing and publishing $\alpha_{t},...,\alpha_{n}$ allows holder $t$ to holder $n$ to obtain $P(t),...,P(n)$. To conclude, the secret share $s_{i}$ gives each secret holder a distinct evaluation on $P$. For holder $1$ to holder $t-1$, $s_{i}$ equals to $P(i)$; for holder $t$ to holder $n$, $s_{i}$ is the key to decrypt the ciphertext $\alpha_{i}$ to obtain the plaintext $P(i)$.
@@
The reveal-verifiability of the recovered secret is provided by the verifiability of each individual secret share from bilinear pairing evaluation. Given that the secret shares are verifiably correct and the $\alpha$s are public, anyone can use the same inputs on Lagrange interpolation to reproduce $k$.
Note that this protocol is also a traditionally defined verifiable secret sharing in the sense that clients cannot communicate incorrect secret shares to secret holders, as the secret shares are all derived deterministically from the same value $g_{1}^{r}$.
@@ V-C Applying to Timed Release Cryptography
Combining all components in the design, we go over the entire process of the proposed cryptographic protocol when it is applied to the timed-release cryptography system. Suppose the client has a message $m$ to encrypt for a period of time; through the published smart contract that stores information of the system, the client knows the public keys of the secret holders $pk_{1},...,pk_{n}$.
@@
    When receiving a request from a client, get the secret share $s_{i}=pk_{i}^{r}=g_{1}^{r\cdot sk_{i}}$.
@@ VI-A Attack Model
Three types of adversaries are taken into account in the attack model, external adversaries, malicious secret holders, and malicious clients. External adversaries are malicious parties that aim to obtain confidential information using public knowledge, specifically, the encrypted messages before decryption time.
@@
Malicious clients also have the potential to disrupt the system if they can compromise the verifiability of decrypted messages. The system fails if clients can falsely accuse honest secret holders of submitting incorrect shares. In this case, honest secret holders are punished and disqualified even if they follow the protocol. Therefore, the proposed protocol must ensure verifiability of secret shares in the existence of malicious clients.
@@ VI-B Message Secrecy Before Decryption
For all generated and secure parameters of holders’ secret keys $sk$ and the clients’ choice of $r$, an external adversary can never use the public knowledge $(g_{1},g_{2},pk_{1},pk_{2},pk_{3},g_{1}^{r},g_{2}^{r},\alpha_{3})$ to derive a secret share before the corresponding holders submit it.
@@
With this base case of three secret holders. We extend this proof to show that message secrecy holds for any size of secret holder.
@@
We further provide proof that the secrecy of the message key $k$ holds in the present of malicious secret holders below the threshold.
@@
Suppose the collaborating malicious secret holders can obtain $k$ before honest holders submit shares, they must be able to gain one extra secret share belonging to the honest holders. That is, they are able to get the secret share using the public knowledge $g_{1},g_{2},g_{1}^{r},g_{2}^{r},pk_{1},...,pk_{n},\alpha_{t+1},...,\alpha_{n}$ plus their own secret keys. Since their own secret keys are randomly generated values, which are irrelevant to any other honest holders’ secret share, they must be able to obtain the secret share using public knowledge. However, lemma VI.2 shows that it is not possible to obtain a secret share using public knowledge, which contradicts the initial assumption and proves the lemma. ∎
@@ VII Performance Evaluation
The most fundamental requirement for a timed-release cryptography system is to decrypt messages at the time users specify. Therefore, decryption time deviation, computed as the difference between the expected decryption time and actual decryption time is measured in this section under various scenarios. Moreover, we provide an estimation of the energy consumption of the proposed system, showing its advantage in energy efficiency.
@@ VII-A Decryption Time Stability
As indicated in Figure 7, the time lock puzzle approach exhibits unreliable decryption time. Its performance deteriorates with increasing encryption duration; for instance, a one-week encryption period results in an average decryption time deviation exceeding two hours. The result indicates two disadvantages of time lock puzzle systems. First, the long error bars illustrate the unstable puzzle-solving time. Second, the average lines at the early decryption area far from zero deviation show that the puzzle difficulty cannot be adjusted perfectly to achieve the desired average solving time. In the case of the Bitcoin blockchain, the hash rate of the network is in a strong increasing trend over time, while the increase of the puzzle difficulty does not strictly follow, resulting in an average block time shorter than expected, thus an accumulated deviation as the encryption duration increases.
@@ VII-B Scalability
In the case of the proposed system, the latency in the decryption process comprises two elements: the time needed for holders to submit secret shares to the blockchain and the time to verify a sufficient number of secret shares. To understand the influence of scaling the number of secret holders on decryption latency, experiments of 3, 10, 20, 30, and 40 secret holders are conducted to process dummy time lock messages.
@@
In Figure 8, the publishing time of the secret shares shows the latency of publishing a sufficient number of shares on the blockchain; the local verification time shows the latency of a secret holder getting enough secret shares verified. As holders process requests concurrently, increasing the number of holders does not extend the time required for the blockchain to accumulate a sufficient number of secret shares. However, as it requires more secret shares to validate a message when the number of secret holders scales, the time required for the secret holders to locally ensure the correct reconstruction of messages increases linearly.
@@ VIII Resilience to Malicious Voting
The sincere and malicious population of voters: In every simulation, a certain population of voters turn into malicious and change their votes based on the winners of the remaining population. So if $l$of $\mathcal{N}$ voters are sincere, $(1-l)$of $\mathcal{N}$ voters change their votes to jeopardize the decision outcomes of the $l$of the voters. We have a total of 100 simulations, where $l$ is varied from 1 to 100, with an increment of 1. In every simulation, we have 100 iterations, wherein the set of malicious voters ($(100-l)$%) changes due to random sampling.
@@
We consider real-world datasets for electing government leaders across different cities or countries (such as UK labor elections, elections in Dublin, Ireland, Oakland and Minneapolis elections in the United States [69, 70, 71, 72], where voters provide either complete or incomplete strict preferential order for the candidates. The election data are taken from the Preflib library [73], which hosts benchmarks for real-world voting and preference data.
We select the datasets to have a mixture of elections that have different numbers of candidates (varying from 5 to 9) or voters (varying from 266 to 36655). As demonstrated in Figure 9, all four elections suffer from a change of winner in the presence of malicious voters. For example, the UK Labour Leadership Election, a small-scale election with only 266 voters, shows a change of winner in the present of 40 strategic voters. While larger-scale elections tolerate more malicious voters, a change of winner is still possible with the collaboration of potential malicious voters with similar preferences.
@@
In summary, a model is formulated to define malicious voting and apply it to real-world election datasets. This model is tested and validated on real-world datasets. We observe that if the proposed system is used, the probability of voters accessing or attacking the votes is negligible. Hence the proposed method is a secure solution to preserve fairness of real-world election outcomes.